Problem 17-40 (Algo) Prorating Overhead Costs (LO 17-1)

Parkeville Company manufactures a single product and started the year with no inventories. Selected information about results for the period just ended include the following:

 






Actual fixed manufacturing overhead$174,000

Actual variable manufacturing overhead
135,000

Applied fixed manufacturing overhead
200,000

Applied variable manufacturing overhead
132,000

Production volume variance
16,000F
Variable overhead efficiency variance
7,000F

 

Eight percent of this period's production has not been sold. There are never any work-in-process inventories.

 

Required:

a. Assume Parkeville writes off all variances to Cost of Goods Sold. Prepare the entries the company would make to record and close out the variances.

b. Assume Parkeville prorates all variances to the appropriate accounts. Prepare the entries the company would make to record and close out the variances.
Answer #1

A)

NoEventGeneral JournalDebitCredit
A1135,000selected answer correctnot attempted


not attempted135,000selected answer correct





B2132,000selected answer correctnot attempted


not attempted132,000selected answer correct





C3132,000selected answer correctnot attempted


10,000selected answer correctnot attempted


not attempted7,000selected answer correct


not attempted135,000selected answer correct





D4200,000selected answer correctnot attempted


not attempted10,000selected answer correct


not attempted16,000selected answer correct


not attempted174,000selected answer correct





E510,000selected answer correctnot attempted


16,000selected answer correctnot attempted


7,000selected answer correctnot attempted


not attempted10,000selected answer correct


not attempted23,000selected answer correct


step 1) for VOH 

132,000- 7,000 F= 125,000

135,000- 125,000= 10,000 U 

135,000- 132,000= 3,000 U


step 2) for FOH

174,000- 200,000= 26,000 F         

200,000- 16,000 F= 184,000

174,000- 184,000= 10,000 F


for entry E COGS: 16,000 F+ 7,000F= 23,000


For B) all journal entries same for entries A-D in part A except E

Fixed overhead price variance                           10,000

Fixed overhead production volume variance      16,000

Variable overhead efficiency variance                   7,000

     Cost of goods sold                                                                   21,160

     Finished goods inventory                                                            1, 840

     Variable overhead spending variance                                      10,000

23,000* 95%(100- 8)= 21,160

23,000* 8%= 1,840

selected answer correctnoselected ansselected answer correct